# Training-free Task-oriented Grasp Generation

### v1

@@ IV-F Failure Modes
Clustering Method Does Not Provide Sufficiently Diverse Clusters: The clustering method used in our approach sometimes fails to generate clusters that are diverse enough. As shown in Fig. 4, many grasps from different clusters are still concentrated in similar regions of the object, limiting the effectiveness of the diversity-based selection process.
@@
Ambiguous Query Images: When the query image is captured from certain viewpoints, some objects or parts of objects in the image may appear ambiguous. For example, as illustrated in Fig. 5, when tasking VLM to select a grasp that grips the lid of the given bottle, it is hard to tell from the image whether the candidate grasps are gripping the lid or the body. This ambiguity can lead to suboptimal grasp suggestions by the VLM.
Limitations of the VLM: The VLM itself has inherent limitations, such as difficulty in understanding complex spatial relationships in 3D scenes or accurately reasoning about physical constraints. Additionally, its responses can sometimes be inconsistent or overly reliant on the specific wording of the task description, which may impact the quality of the selected grasp.
Addressing these failure modes will require improvements in clustering techniques, better handling of viewpoint selection for query images, and leveraging more advanced VLMs with enhanced spatial reasoning and contextual understanding capabilities.
V Conclusion
In this paper, we introduced a training-free pipeline for task-oriented grasp generation that integrates pre-trained grasp generation models with vision-language models (VLMs) to select grasps aligned with specific task requirements. Our approach demonstrated significant improvements over a baseline system, achieving up to a 36.9% absolute gain in overall success rate. Among the evaluated strategies, the Contact Point Generation (CPG) method showed the best task compliance, while other methods exhibited complementary strengths in grasp success and task-specific alignment. Despite its effectiveness, we identified limitations such as insufficient grasp diversity, query image ambiguities, and VLM constraints, providing directions for future improvements. This study highlights the potential of combining pre-trained grasp models and VLMs for flexible, efficient task-oriented manipulation without additional training, offering a promising framework for advancing robotic grasping capabilities.

### v2

@@ IV-D Failure Modes
Clustering Method Does Not Provide Sufficiently Diverse Clusters: The clustering method used in our approach sometimes fails to generate clusters that are diverse enough. As shown in Figure 5 (a), many grasps from different clusters are still concentrated in similar regions of the object, limiting the effectiveness of the diversity-based selection process.
@@
Ambiguous Query Images: When the query image is captured from certain viewpoints, some objects or parts of objects in the image may appear ambiguous. For example, as illustrated in Figure 5 (b), when tasking VLM to select a grasp that grips the lid of the given bottle, it is hard to tell from the image whether the candidate grasps are gripping the lid or the body. This ambiguity can lead to suboptimal grasp suggestions by the VLM.
VLM Limitations: The VLM itself has inherent constraints, including challenges in reasoning about complex spatial relationships and physical constraints in 3D scenes [25]. As shown in Figure 5(c), the VLM selects a grasp that results in instability, while in Figure 5(d), the contact point specified by the VLM is completely misaligned with the object.
V Real-world Experiment
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: The top row illustrates the task ”use the brush,” while the bottom row corresponds to the task ”pass me the brush.” The robot selects different grasping points based on the task description, demonstrating its ability to adapt grasp strategies according to task-oriented requirements.
To test if our proposed method works in the real world, we extends the evaluation of our proposed method beyond simulation environments to assess its performance in real-world robotic manipulation tasks.
V-A Experimental Setup
Our hardware configuration employed a Franka Emika Panda robotic arm equipped with an Intel RealSense L515 RGB-D camera mounted on the end-effector link (eye-in-hand). We use the RGB and depth image to generate the point clouds. We use a single singe image to query the VLM and generate the grasps, thus mimic our setup in the simulator.
We utilized Contact GraspNet (CGN) [3] to generate unconditional grasps for each object and used Gemini-2.0-Pro [33] as the VLM model. We use the CPG (Contact Point Generation) strategy as it performs the best III, in which we directly ask the VLM to generate the best grasp point that satisfying the task description. Interesting, we find that the VLM can correctly select the corresponding grasps for the target object, without using a object detection and segmentation pipeline as in the common setup, indicating the use of VLM to further simplify the grasp generation pipeline.
We evaluated grasp performance on 10 objects across three different scenes (Figure 7):
  • Common Household Items: The robot interacted with everyday objects, such as an umbrella, grey lotion, and a grey brush, arranged in cluttered settings.
  • Kitchen Utensils: The robot handled various kitchen items, including a red pot, black spoon, red cup, whisk, and muddler, positioned among multiple objects.
  • Hand Tools: The robot operated tools such as an axe and a screwdriver in densely arranged scenes.
For each object, we evaluated performance on two types of tasks:
  • Use the Tool: Success is measured by achieving a stable grasp appropriate for the intended use. For example, the robot should securely hold the red pot by its handle for cooking or firmly grip the whisk for mixing.
  • Handover: Success is defined by passing the object to a person in a way that allows for easy and safe grasping. For instance, the robot should pass a hammer by grasping its head so that the user can easily take hold of the handle.
Each object-task combination was evaluated over three trials.
Metric Household Kitchen Tools
Grasp Success 0.72 0.90 0.83
Task Compliance 0.94 0.80 0.83
Combined Success 0.68 0.70 0.75
TABLE IV: Real-world experiment results
(a) Common Item
(b) Kitchen
(c) Hand Tools
Figure 7: Three real-world scenes
V-B Result
Table IV presents the results of our real-world experiment. All three scenes consist of multiple objects, requiring the system to generate collision-free grasps. By integrating ContactNet’s [3] capability to generate collision-free grasps from partial point clouds with the VLM’s ability to predict grasping points based on the task description, the TOG system remains both simple and effective. Requiring only a single pair of RGB and depth images as input, it achieves an overall success rate of 70%.
We observe that the VLM generally achieves a high success rate in indicating reasonable grasp points. However, its performance declines when ambiguity arises due to the type of object and the viewpoint. For example, when viewed from a top-down perspective, the muddler’s handle is difficult to distinguish, making it challenging to determine the correct grasping location. This aligns with our observations in the simulated experiment (Section IV-D).
VI Conclusion
In this paper, we introduced TOG, a training-free system for task-oriented grasp generation that integrates pre-trained grasp generation models with vision-language models (VLMs) to select grasps aligned with specific task requirements. We evaluated the proposed system in both simulated and real-world environments, demonstrating its effectiveness across various task setups.
Crucially, we find that when leveraging a powerful VLM, simply prompting the model to indicate where to grasp and selecting from a set of feasible grasps generated by a pre-trained grasp generation model — the CPG strategy — yields the best performance.
Despite its effectiveness, we identified several limitations, including insufficient grasp diversity, query image ambiguities, and constraints inherent to VLMs, highlighting areas for future improvement. This study underscores the potential of combining pre-trained grasp models with VLMs to enable flexible, efficient task-oriented manipulation without additional training, offering a promising framework for advancing robotic grasping capabilities.